OpenStudy (aonz):

Integration rotation bounded by 2 curves question :) Help Please

OpenStudy (aonz):

(a)On the same set of axis sketch the curves y = 18-2x^2 and \[\huge y =\sqrt{9-x ^{2}}\] (b) Find the area bounded by the 2 curves (c)Find the volume of the solid formed when this region is rotated about the x axis.